OpenStudy (anonymous):

Help pleaseeee.. The function f(x) = 12(4)x represents the growth of a fly population every year in a remote swamp. Jackie wants to manipulate the formula to an equivalent form that calculates three times a year, not just once a year. Which function is correct for Jackie's purpose, and what is the new growth rate?

OpenStudy (anonymous):

f(x) = 4(4)x; growth rate 400% f(x) = 4(4)3x, growth rate 4% f(x) = 4(1.59)3x; growth rate 59% f(x) = 4(1.59)x; growth rate 4% These are my options

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

Can you please help?

OpenStudy (anonymous):

@Luigi0210 @KingGeorge

OpenStudy (anonymous):

@goformit100